Three unknown numbers are given. The second number is one more than the first. The third number is the sum of the first and second plus one. If the sum of the first and third is 21, what is the second number?
Answer by stanbon(75887)

Three unknown numbers are given::: x,y,z
---------------------------------------------
The second number is one more than the first.
y = x+1
---------------------------------------------
The third number is the sum of the first and second plus one.
z = x+y+1
---------------------------------------------
If the sum of the first and third is 21
x+z = 21
---------------------
what is the second number?
================================
Rearrange the 3 equations:
x + 0y + z = 21
x + y - z = -1
x + -y +0z = -1
-----------------------
I used a matrix method to get:
x = 6 1/3
y = 7 1/3
z = 14 2/3
================
Cheers,
Stan H.
================